# One-Embedding-Fits-All: Efficient Zero-Shot Time Series Forecasting by a Model Zoo

### v1

@@ Appendix B Dataset in Gift-Eval Benchmark
The GIFT-Eval benchmark GE is a large-scale, general-purpose evaluation suite designed specifically for testing the zero-shot and universal forecasting capabilities of time series foundation models (TSFMs). The full list of datasets and their metadata is provided in Table 6. GIFT-Eval comprises 23 real-world datasets with 97 configurations collected from diverse sources, encompassing over 144,000 time series and 177 million individual observations. Each dataset is labeled by key structural features, including domain, prediction length, frequency, and number of variates, enabling fine-grained group-wise performance analysis along the following axes:
@@
Appendix C Implementation Details
@@
C.1 Co-embedding Extractor Details
@@
$\mathcal{L}_{\text{Reconstruction}}$ $=\min_{\Theta}\;\|x_{\phi_{m}}-\hat{x}_{\phi_{m}}\|_{2}^{2},$
$\mathcal{L}_{\text{Constraint}}$ $=-\sum\log\frac{\exp(\text{sim}(s,s^{\prime}))}{\sum\exp(\text{sim}(s,s^{\prime\prime}))},$
$\mathcal{L}_{\text{Transfer}}$ $=\sum_{\mathbf{x}_{i},\mathbf{x}_{j}\in\mathcal{D}}\left\|g_{i,j}-\text{sim}(E(\mathbf{x}_{i}),E(\mathbf{x}_{j}))\right\|_{2}^{2},$
where $x_{\phi_{m}}$ denotes the input time series, $\hat{x}_{\phi_{m}}$ is its reconstruction, and $\text{sim}(\cdot,\cdot)$ computes cosine similarity. Our transferability score is designed using the $1$-MSE formulation, defined as $g_{i,j}=1-\text{MSE}(\phi_{i},\mathcal{D}_{j})$, where $\phi_{i}$ is the foundation model pretrained on dataset $\mathcal{D}_{i}$, and $\text{MSE}(\phi_{i},\mathcal{D}_{j})$ measures the forecasting error of $\phi_{i}$ on dataset $\mathcal{D}_{j}$. This design quantifies cross-task transferability and aligns the learned embedding space with actual generalization ability across datasets.
@@
Approximately 300,000 subsequences are sampled from M3, M4, and Tourism datasets to compute cross-task transferability scores and to construct advantage subsets for unseen tasks. In contrast, Moirai pretrains on LOTSA (2.7B points) and Sundial on TimeBench (1T points), highlighting that our extractor relies on relatively modest-scale data with short subsequences (length 36) to ensure computational efficiency.
C.2 Zero-shot Forecasting Benchmark Details
In the Gift-Eval benchmark, the zoo representation set $\mathcal{D}$ contains $n=1000$ subsequences sampled from the pre-training datasets of TSFMs. We apply the subset thresholding strategy in Equation 3 with an adaptive threshold $\tau=1$, and define the model weight $w_{m}$ as the inverse square root of the number of samples in $\mathcal{D}_{m}^{\text{adv}}$ that exceed the threshold. The co-embedding extractor $\psi$ maps each $\mathbb{R}^{36}$ input segment into a $\mathbb{R}^{128}$ latent representation, resulting in the model zoo representation library $\mathbf{R}{\text{zoo}}\in\mathbb{R}^{M\times 128}$. For task representations, we process the multivariate time series channel-wise: from each channel, five random $\mathbb{R}^{36}$ segments are sampled, encoded with $\psi$, and then averaged to obtain channel embeddings $\mu_{c}\in\mathbb{R}^{128}$. Stacking across all channels yields the task embedding matrix $\mu=[\mu_{1},\mu_{2},\dots,\mu_{C}]^{\top}\in\mathbb{R}^{C\times 128}.$ Model rankings are derived by computing weighted cosine similarities between $\mu$ and $\mathbf{R}_{\text{zoo}}$.
  • Chronos chronos : We follow the official evaluation setup using num_samples = 20. In the GIFT-Eval benchmark, we use four upgraded versions released in November 2024: chronos-bolt-tiny, chronos-bolt-mini, chronos-bolt-small, and chronos-bolt-base, denoted respectively as Chr.bT, Chr.bM, Chr.bS, and Chr.bB.
  • Moirai MOIRAI : We follow the official evaluation setup using num_samples = 100 and patch_size = 32. In the GIFT-Eval, we set context_length = 4000, and feat_dynamic_real_dim is set dynamically based on the dataset. We use three different model sizes released in March 2024: moirai-1.1-R-small, moirai-1.1-R-base, and moirai-1.1-R-large, denoted as Moi.S, Moi.B, and Moi.L, respectively.
  • TimesFM TimesFM : We set context_length = 512 in the GIFT-Eval. We use two different model sizes: timesfm-1.0-200m released in March 2024 and timesfm-2.0-500m released in December 2024, denoted as TFM.1 and TFM.2, respectively.
  • VisionTS VisionTS : Due to the limitations of VisionTS, the maximum context length is 96, while the minimum length depends on the shortest input supported by each dataset. We follow the official implementation and include three model sizes released in August 2024: mae_visualize_vit_base, mae_visualize_vit_large, and mae_visualize_vit_huge, denoted as Vis.B, Vis.L, and Vis.H, respectively.
  • Sundial liu2025sundialfamilyhighlycapable : We follow the official evaluation setup using num_samples = 100 and context_length = 512 in the GIFT-Eval. We use the only public model size: Sundial_base_128m released in June 2025, denoted as Sun.B.
Model selection baseline. We adapted LogME LogME , originally designed for CV and NLP tasks, to the time-series forecasting setting by directly applying its core principle of measuring model–data fit. Concretely, we computed LogME scores for each TSFM–task pair and used these scores to generate a model ranking, replacing ZooCast’s ranking module. However, this approach faces inherent limitations: it cannot effectively handle multivariate time-series tasks, so our evaluation was restricted to 32 univariate datasets in the GIFT-Eval benchmark. Moreover, LogME is computationally inefficient, since it requires forward passes of all TSFMs for every new task and does not allow reusing precomputed results, leading to substantial overhead compared with our design.
Appendix D Further Information on ZooCast
@@
D.1 Accuracy: Embedding-Aware Model Weighting
During the precompute stage of ZooCast, we not only construct the model zoo representation library $\mathbf{R}_{\text{zoo}}$, but also record a weighting factor $w_{m}$ for each model $\phi_{m}$ that captures the relative strength of each model. This design is motivated by a key observation: although state-of-the-art (SOTA) models continue to evolve, at any given time, some models significantly outperform others across the majority of tasks. Treating all models equally in the selection phase would dilute this prior knowledge and underutilize the generalizability of strong models.
But how can we determine which models in the zoo are superior when the downstream task is unknown? We determine model weights during the construction of the advantage subset $\mathcal{D}_{m}^{\text{adv}}$ as described in Equation 3 of the main paper, where $\tau$ is an adaptive threshold. The number of samples that satisfy the threshold condition for model $\phi_{m}$ is denoted by $d_{m}$:
$d_{m}=\left|\left\{\mathbf{x}_{i}\mid\mathbb{E}_{\phi_{k}\neq\phi_{m}}[\mathcal{L}(\phi_{k}(\mathbf{x}_{i}))]-\mathcal{L}(\phi_{m}(\mathbf{x}_{i}))>\tau\right\}\right|,\quad w_{m}\;=\;\frac{1}{\sqrt{d_{m}}}$
@@ D.1 Accuracy: Embedding-Aware Model Weighting
During selection, these weights are used to modulate similarity scores between task embeddings $\mu_{c}$ and model representations $\mathbf{r}_{m}$:
$\text{sim}_{m,c}\;=\;w_{m}\cdot\frac{\mathbf{r}_{m}^{\top}\mu_{c}}{\|\mathbf{r}_{m}\|\,\|\mu_{c}\|}.$
This formulation favors models that demonstrated stronger performance during the precompute phase, even when their raw embedding similarity is slightly lower. The final similarity matrix $\mathbf{A}=[\text{dist}_{m,c}]_{C\times M}$ is used for ranking and model selection.
D.2 Scalability: Seamless Model Zoo Expansion
ZooCast enables rapid and scalable integration of new models through an efficient update mechanism that avoids redundant computation. Assume the current model zoo consists of $M$ models, for which the full evaluation matrix $\mathbf{E}^{\text{old}}\in\mathbb{R}^{M\times n}$ has been precomputed, where $n$ denotes the number of time series samples in the fixed zoo representation set $\mathcal{D}$. Correspondingly, the model zoo representation library is maintained as $\mathbf{R}_{\text{zoo}}^{\text{old}}\in\mathbb{R}^{M\times D}$, where $D$ is the dimensionality of the representation vector for each semantic unit.
@@ D.2 Scalability: Seamless Model Zoo Expansion
$\mathbf{E}^{\text{old}}$ $\in\mathbb{R}^{M\times n}\quad\xrightarrow{\text{\makebox[85.35826pt]{Add $\phi_{M+1}$ results}}}\quad\mathbf{E}^{\text{new}}\in\mathbb{R}^{(M+1)\times n}\quad\text{(< 1 min)}$
$\mathbf{R}_{\text{zoo}}^{\text{old}}$ $\in\mathbb{R}^{M\times D}\quad\xrightarrow{\text{\makebox[85.35826pt]{Append new repr.}}}\quad\mathbf{R}_{\text{zoo}}^{\text{new}}\in\mathbb{R}^{(M+1)\times D}\quad\text{(< 1 s)}$
@@
D.3 Efficiency: Minimal Computational Complexity and Memory Usage
@@
Repr-based (ZooCast). Our method introduces a precompute-and-select paradigm that fundamentally changes the scaling behavior of model zoo inference. In the precompute phase, each of the $M$ models is forwarded on a small subset of $n$ samples to build its representation, yielding a one-time cost of $\mathcal{O}(Mn)$. This cost is paid only once when the zoo is constructed, and never repeated for subsequent tasks. When averaged over $U$ downstream tasks, the per-task amortized cost is $\mathcal{O}({Mn}/{U})$, which converges to $\mathcal{O}(1)$ as $U$ grows large with fixed $n\ll N$. In the selection phase, instead of re-running models, ZooCast only performs cosine similarity between the task embedding $\mu\in\mathbb{R}^{C\times D}$ and the precomputed model embeddings $\mathbf{R}_{\text{zoo}}\in\mathbb{R}^{M\times D}$. These are low-dimensional ($D\leq 128$) vector operations, so all computations reduce to matrix slicing, dot products, and max-pooling. Each of these costs constant time per model, and with parallelization, the effective cost per task is $\mathcal{O}(1)$. This sharp contrast with forward-based approaches highlights the scalability of ZooCast: regardless of the zoo size $M$, the selection stage does not grow in complexity. Finally, in the forecast phase, ZooCast generates predictions only with the top-selected models, requiring $\mathcal{O}(N)$ operations per task. Thus, the entire pipeline reduces the total per-task cost from $\mathcal{O}(Mn+N)$ in forward-based methods to $\mathcal{O}({Mn}/{U}+1+N)$, which approaches $\mathcal{O}(N)$ for large $U$.
Memory usage (ZooCast). ZooCast’s memory footprint is dominated by (1) the precomputed error matrix $\mathbf{E}\in\mathbb{R}^{M\times n}$ and (2) the zoo representation library $\mathbf{R}_{\text{zoo}}\in\mathbb{R}^{M\times D}$. Both scale linearly with $M$. In practice, at $M=10$, storage is only about 0.2 MB, and even at $M=1000$, storage remains manageable (around 20 MB).

### v2

@@ Appendix B Dataset in Gift-Eval Benchmark
The GIFT-Eval benchmark (Aksu et al., 2024) is a large-scale, general-purpose evaluation suite designed specifically for testing the zero-shot and universal forecasting capabilities of time series foundation models (TSFMs). The full list of datasets and their metadata is provided in Table 6. GIFT-Eval comprises 23 real-world datasets with 97 configurations collected from diverse sources, encompassing over 144,000 time series and 177 million individual observations. Each dataset is labeled by key structural features, including domain, prediction length, frequency, and number of variates, enabling fine-grained group-wise performance analysis along the following axes:
@@
Appendix C Extended Related Work
We provide additional details on model selection methods relevant to our work.
Model Selection Methods. Early model selection in transfer learning relied on brute-force fine-tuning, exhaustively adapting each model to the target task. While accurate, this was infeasible at scale. To reduce cost, forward-based methods emerged, approximating model suitability through lightweight inference. LEEP (Nguyen et al., 2020) estimates the conditional distribution between source and target labels via a single forward pass, while LogME (You et al., 2021) computes a marginal likelihood to evaluate feature–label alignment for both classification and regression, though it assumes Gaussian labels and is costly in high dimensions. Despite efficiency gains, forward-based methods still depend on labeled target data and light training, making them incompatible with strict zero-shot settings. To eliminate the need for target labels or forward passes, representation-based methods compare tasks in a shared embedding space. TASK2VEC (Achille et al., 2019) pioneered this idea using Fisher information but is tied to classification architectures. TS2Vec (Yue et al., 2022) introduced multi-scale temporal embeddings for time series, but its data-driven embeddings are not aligned with model behavior. SimMTM (Dong et al., 2023) adopted masked modeling to predict missing values, which is lightweight and effective for pretraining but lacks mechanisms for transferability or model selection. TimesURL (Liu & Chen, 2024) enhanced TS2Vec with frequency-time augmentations, yet remains computationally heavy and requires post hoc model ranking. ModelSpider (Zhang et al., 2023) framed model selection as a learning-to-rank task, effective in classification but requiring expensive fine-tuned labels and not extended to regression or time series forecasting. Overall, these methods remain limited by structural rigidity, task specificity, computational overhead, and lack of integrated ranking mechanisms, leaving time series zero-shot selection unresolved.
Appendix D Zero-shot Forecasting Benchmark Details
@@
In the Gift-Eval benchmark, the zoo representation set $\mathcal{D}$ contains $n=1000$ subsequences sampled from the pre-training datasets of TSFMs. We apply the subset thresholding strategy in Equation 6 with an adaptive threshold $\tau=1$, and define the model weight $w_{m}$ as the inverse square root of the number of samples in $\mathcal{D}_{m}^{\text{adv}}$ that exceed the threshold. The co-embedding extractor $\psi$ maps each $\mathbb{R}^{36}$ input segment into a $\mathbb{R}^{128}$ latent representation, resulting in the model zoo representation library ${\bm{R}}_{\text{zoo}}\in\mathbb{R}^{M\times 128}$. For task representations, we process the multivariate time series channel-wise: from each channel, five random $\mathbb{R}^{36}$ segments are sampled, encoded with $\psi$, and then averaged to obtain channel embeddings ${\bm{\mu}}_{c}\in\mathbb{R}^{128}$. Stacking across all channels yields the task embedding matrix ${\bm{R}}_{\text{task}}=[\mu_{1},\mu_{2},\dots,{\bm{\mu}}_{C}]^{\top}\in\mathbb{R}^{C\times 128}.$ Model rankings are derived by computing weighted cosine similarities between $\mu$ and ${\bm{R}}_{\text{zoo}}$.
  • Chronos (Ansari et al., 2024): We follow the official evaluation setup using num_samples = 20. In the GIFT-Eval benchmark, we use four upgraded versions released in November 2024: chronos-bolt-tiny, chronos-bolt-mini, chronos-bolt-small, and chronos-bolt-base, denoted respectively as Chr.bT, Chr.bM, Chr.bS, and Chr.bB.
  • Moirai (Woo et al., 2024): We follow the official evaluation setup using num_samples = 100 and patch_size = 32. In the GIFT-Eval, we set context_length = 4000, and feat_dynamic_real_dim is set dynamically based on the dataset. We use three different model sizes released in March 2024: moirai-1.1-R-small, moirai-1.1-R-base, and moirai-1.1-R-large, denoted as Moi.S, Moi.B, and Moi.L, respectively.
  • TimesFM (Das et al., 2024): We set context_length = 512 in the GIFT-Eval. We use two different model sizes: timesfm-1.0-200m released in March 2024 and timesfm-2.0-500m released in December 2024, denoted as TFM.1 and TFM.2, respectively.
  • VisionTS (Chen et al., 2025): Due to the limitations of VisionTS, the maximum context length is 96, while the minimum length depends on the shortest input supported by each dataset. We follow the official implementation and include three model sizes released in August 2024: mae_visualize_vit_base, mae_visualize_vit_large, and mae_visualize_vit_huge, denoted as Vis.B, Vis.L, and Vis.H, respectively.
  • Sundial (Liu et al., 2025): We follow the official evaluation setup using num_samples = 100 and context_length = 512 in the GIFT-Eval. We use the only public model size: Sundial_base_128m released in June 2025, denoted as Sun.B.
Model selection baseline. We adapted LogME (You et al., 2021), originally designed for CV and NLP tasks, to the time-series forecasting setting by directly applying its core principle of measuring model–data fit. Concretely, we computed LogME scores for each TSFM–task pair and used these scores to generate a model ranking, replacing ZooCast’s ranking module. However, this approach faces inherent limitations: it cannot effectively handle multivariate time-series tasks, so our evaluation was restricted to 32 univariate datasets in the GIFT-Eval benchmark. Moreover, LogME is computationally inefficient, since it requires forward passes of all TSFMs for every new task and does not allow reusing precomputed results, leading to substantial overhead compared with our design.
Appendix E Additional Methodological Details
This section provides the extended technical details that complement the main text. Specifically, we present (i) the formulation of advantage scores for constructing model-specific subsets, (ii) the full design of the co-embedding extractor with its loss functions, training setup, and datasets.
E.1 Advantage Score Computation
We first forward all models on the zoo characterization set $\mathcal{D}$ to obtain the MSE matrix ${\bm{E}}\in\mathbb{R}^{M\times n}$, where $E_{m,i}$ is the MSE of $\phi_{m}$ on subsequence ${\bm{x}}_{i}$. Let the inter-model error variance on sample ${\bm{x}}_{i}$ be $\sigma_{i}\;=\;\operatorname{std}\!\big(\{E_{m,i}\}_{m=1}^{M}\big)$, and let $\bar{\sigma}$ and $\hat{\sigma}$ be the mean and standard deviation of $\{\sigma_{i}\}$. The per-sample advantage score for model $\phi_{m}$ is then defined as
$s_{m,i}\;=\;\left(\frac{1}{M-1}\sum_{k\neq m}E_{k,i}\;-\;E_{m,i}\right)\cdot\frac{\sigma_{i}-\bar{\sigma}}{\hat{\sigma}},\quad\mathcal{D}_{m}^{\text{adv}}\;=\;\left\{\,{\bm{x}}_{i}\in\mathcal{D}\;\middle|\;s_{m,i}\;>\;\tau\,\right\}.$ (6)
This score jointly reflects (i) how much better model $\phi_{m}$ performs compared to its peers on sample ${\bm{x}}_{i}$, and (ii) how discriminative the sample is based on inter-model disagreement. A sample is included in $\mathcal{D}_{m}^{\text{adv}}$ if $s_{m,i}>\tau$, where $\tau$ is an adaptive threshold. This design ensures that advantage subsets focus on high-variance, informative samples that best reveal model-specific strengths. The cardinality $d_{m}\;=\;\big|\mathcal{D}_{m}^{\text{adv}}\big|$ is used to derive the model weight $w_{m}$ in selection (details in Appendix F.1).
E.2 Co-embedding Extractor Details
@@
$\mathcal{L}_{\text{Reconstruction}}$ $=\min_{\Theta}\;\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2},\quad{\bm{x}}\in\mathcal{D}^{*},$
$\mathcal{L}_{\text{Constraint}}$ $=-\sum_{{\bm{x}}\in\mathcal{D}^{*}}\sum_{\tilde{{\bm{x}}}\in\mathcal{M}({\bm{x}})}\log\frac{\exp\!\left(\operatorname{sim}(\psi({\bm{x}}),\psi(\tilde{{\bm{x}}}))\right)}{\sum_{{\bm{x}}^{\prime}\in\mathcal{D}^{*}}\exp\!\left(\operatorname{sim}(\psi({\bm{x}}),\psi({\bm{x}}^{\prime}))\right)},$
$\mathcal{L}_{\text{Transfer}}$ $=\sum_{{\bm{x}}_{i}\in\mathcal{D}_{i},\;{\bm{x}}_{j}\in\mathcal{D}_{j}}\left\|g_{i,j}-\operatorname{sim}\!\big(\psi({\bm{x}}_{i}),\psi({\bm{x}}_{j})\big)\right\|_{2}^{2}.$
[…]
@@
Approximately 300,000 subsequences are sampled from M3, M5, and Tourism datasets to compute cross-task transferability scores and to construct advantage subsets for unseen tasks. In contrast, Moirai pretrains on LOTSA (2.7B points) and Sundial on TimeBench (1T points), highlighting that our extractor relies on relatively modest-scale data with short subsequences (length 36) to ensure computational efficiency.
Appendix F Further Information on ZooCast
@@
F.1 Accuracy: Embedding-Aware Model Weighting
During the precompute stage of ZooCast, we not only construct the model zoo representation library ${\bm{R}}_{\text{zoo}}$, but also record a weighting factor $w_{m}$ for each model $\phi_{m}$ that reflects its relative strength. This design is motivated by a key observation: although state-of-the-art (SOTA) models continue to evolve, at any given time, some models significantly outperform others across the majority of tasks. Treating all models equally in the selection phase would dilute this prior knowledge and underutilize the generalizability of strong models.
But how can we determine which models in the zoo are superior when the downstream task is unknown? We determine model weights during the construction of the advantage subset $\mathcal{D}_{m}^{\text{adv}}$ as described in Appendix E.1, where $\tau$ is an adaptive threshold. The number of samples that satisfy the threshold condition for model $\phi_{m}$ is denoted by $d_{m}$:
$d_{m}\;=\;\big|\mathcal{D}_{m}^{\text{adv}}\big|,\quad w_{m}\;=\;\frac{1}{\sqrt{d_{m}}}$
@@ F.1 Accuracy: Embedding-Aware Model Weighting
During selection, these weights are used to modulate similarity scores (equation 5), which favors models that demonstrated stronger performance during the precompute phase, even when their raw embedding similarity is slightly lower.
F.2 Scalability: Seamless Model Zoo Expansion
ZooCast enables rapid and scalable integration of new models through an efficient update mechanism that avoids redundant computation. Assume the current model zoo consists of $M$ models, for which the full evaluation matrix ${\bm{E}}^{\text{old}}\in\mathbb{R}^{M\times n}$ has been precomputed (provided in Appendix E.1), where $n$ denotes the number of time series samples in the fixed zoo representation set $\mathcal{D}$. Correspondingly, the model zoo representation library is maintained as ${\bm{R}}_{\text{zoo}}^{\text{old}}\in\mathbb{R}^{M\times D}$, where $D$ is the dimensionality of the representation vector for each semantic unit.
@@ F.2 Scalability: Seamless Model Zoo Expansion
${\bm{E}}^{\text{old}}$ $\in\mathbb{R}^{M\times n}\quad\xrightarrow{\text{\makebox[85.35826pt]{Add $\phi_{M+1}$ results}}}\quad{\bm{E}}^{\text{new}}\in\mathbb{R}^{(M+1)\times n}\quad\text{(< 1 min)}$
${\bm{R}}_{\text{zoo}}^{\text{old}}$ $\in\mathbb{R}^{M\times D}\quad\xrightarrow{\text{\makebox[85.35826pt]{Append new repr.}}}\quad{\bm{R}}_{\text{zoo}}^{\text{new}}\in\mathbb{R}^{(M+1)\times D}\quad\text{(< 1 s)}$
@@
F.3 Efficiency: Minimal Computational Complexity and Memory Usage
@@
Repr-based (ZooCast). Our method introduces a precompute-and-select paradigm that fundamentally changes the scaling behavior of model zoo inference. In the precompute phase, each of the $M$ models is forwarded on a small subset of $n$ samples to build its representation, yielding a one-time cost of $\mathcal{O}(Mn)$. This cost is paid only once when the zoo is constructed, and never repeated for subsequent tasks. When averaged over $U$ downstream tasks, the per-task amortized cost is $\mathcal{O}({Mn}/{U})$, which converges to $\mathcal{O}(1)$ as $U$ grows large with fixed $n\ll N$. In the selection phase, instead of re-running models, ZooCast only performs cosine similarity between the task representation ${\bm{R}}_{\text{task}}\in\mathbb{R}^{C\times D}$ and the precomputed model zoo embeddings ${\bm{R}}_{\text{zoo}}\in\mathbb{R}^{M\times D}$.These are low-dimensional ($D\leq 128$) vector operations, so all computations reduce to matrix slicing, dot products, and max-pooling. Each of these costs constant time per model, and with parallelization, the effective cost per task is $\mathcal{O}(1)$. This sharp contrast with forward-based approaches highlights the scalability of ZooCast: regardless of the zoo size $M$, the selection stage does not grow in complexity. Finally, in the forecast phase, ZooCast generates predictions only with the top-selected models, requiring $\mathcal{O}(N)$ operations per task. Thus, the entire pipeline reduces the total per-task cost from $\mathcal{O}(Mn+N)$ in forward-based methods to $\mathcal{O}({Mn}/{U}+1+N)$, which approaches $\mathcal{O}(N)$ for large $U$.
Memory usage (ZooCast). ZooCast’s memory footprint is dominated by (1) the precomputed error matrix ${\bm{E}}\in\mathbb{R}^{M\times n}$ and the zoo representation library ${\bm{R}}_{\text{zoo}}\in\mathbb{R}^{M\times D}$. Both scale linearly with $M$. In practice, at $M=10$, storage is only about 0.2 MB, and even at $M=1000$, storage remains manageable (around 20 MB).
@@
Appendix G Dataset Usage Summary
We summarize the three types of datasets used in this work and clarify collection effort and leakage safeguards.
(1) Auxiliary training set for the co-embedding extractor ($\mathcal{D}^{*}$). We sample approximately 300,000 subsequences (length $36$) from public time-series corpora (M3, M5, Tourism) to train the extractor $\psi$ and to supervise the transferability loss (Appendix E.2). This set is fully disjoint from both TSFM pretraining corpora and downstream evaluation data. Its scale is moderate by design (short segments, standard datasets), which keeps training efficient while providing broad pattern diversity. Once $\psi$ is trained, $\mathcal{D}^{*}$ is not accessed again during the precompute or evaluation phases, preventing any backdoor leakage into selection or forecasting.
(2) Zoo characterization set for advantage subsets and model embeddings ($\mathcal{D}$). For each TSFM $\phi_{m}$ in the zoo, we randomly draw only hundreds of subsequences from the model’s public pretraining pools (as released by the original TSFM papers, e.g., Chronos, Moirai). This small, diverse set is used once to compute the error matrix ${\bm{E}}\in\mathbb{R}^{M\times n}$, construct the advantage subsets $\mathcal{D}_{m}^{\text{adv}}$, and build the model zoo representation library ${\bm{R}}_{\text{zoo}}$ (Appendix E.1). Collection is straightforward (public sources, random sampling), and because $\mathcal{D}$ comes from TSFM pretraining data rather than downstream benchmarks, there is no leakage path into the evaluation sets. After precompute, ${\bm{R}}_{\text{zoo}}$ is fixed and reused across tasks; neither labels nor raw series from $\mathcal{D}$ are needed at test time.
(3) Downstream evaluation data (GIFT-Eval). All zero-shot results are reported on the public GIFT-Eval benchmark (23 datasets, 97 configurations; Appendix B). GIFT-Eval is used only for evaluation and for producing model rankings at test time via lightweight similarity between ${\bm{R}}_{\text{task}}$ and ${\bm{R}}_{\text{zoo}}$; it is never used to train $\psi$ or to construct ${\bm{R}}_{\text{zoo}}$. Thus, evaluation happens under a strict zero-shot protocol with clear phase isolation.
Leakage safeguards and phase isolation. Data usage follows a clean separation: $\mathcal{D}^{*}$ is used exclusively to train $\psi$; $\mathcal{D}$ (public TSFM pretraining pools with only small random samples) is used exclusively to precompute $\bm{E}$}, $\mathcal{D}_{m}^{\text{adv}}$, and ${\bm{R}}_{\text{zoo}}$; GIFT-Eval is reserved exclusively for downstream evaluation. No dataset is reused across phases in a way that could leak target information into training or precompute. Moreover, our method requires diversity but not scale: the precompute step relies on small, randomly sampled and publicly documented sources, which further reduces collection burden and eliminates hidden data dependencies. Together, these choices ensure simple collection, transparency, and no leakage risk across all stages.
Appendix H Limitations
While ZooCast introduces a novel and efficient framework for zero-shot model selection in time-series forecasting, several limitations remain. The paradigm of using a model zoo for prediction, though increasingly common in vision and language domains, is still nascent in the time-series community. Consequently, wider adoption of such frameworks may require additional time for ecosystem development, tooling, interpretability improvements, and community validation. Furthermore, the effectiveness of our method currently depends on a curated zoo and a fixed representation extractor trained on pre-defined datasets, which may not fully capture the diversity of future task distributions. In addition, the model selection process itself is somewhat constrained: the extractor only supports fixed-length inputs (length 36) and does not yet accommodate dynamic sequence lengths or multimodal/co-variable information from domains such as text or images. The integration scheme employed in our current design is also relatively simplistic, discarding potentially valuable multi-quantile prediction information and lacking mechanisms for dynamic task adaptation, thereby limiting ensemble expressiveness. Finally, the performance ceiling of ZooCast is inherently bounded by the strongest individual TSFM in the zoo; when integration is too basic or ensembles are not employed, the method cannot consistently surpass the best single model. These limitations highlight important avenues for future work, including more flexible representation learning, richer integration strategies, and adaptive ensembles that can better exploit the strengths of diverse TSFMs.
Appendix I The Use of Large Language Models
In preparing this manuscript, we employed large language models (LLMs) primarily as auxiliary tools for non-scientific tasks, including language polishing, table formatting, and mathematical notation checking. These tools were used to improve clarity and presentation consistency without influencing the research design, experimental results, or scientific conclusions. All technical content, algorithms, and experiments remain the original contributions of the authors.